How do you solve by elimination using multiplication:
-3x+4y=12
2x+y=-8

Answer by unlockmath(1688)

Hello,
Let's multiply the second equation by -4 which will give us:

-3x+4y=12
-8x-4y=32
Now we can add both equations to give us:
-11x=44
Now we can divide both sides by -11 to give us:
x=-4
Plug in -4 to the original equation to find y.
-3(-4)+4y=12
12+4y=12 Subtract 12 to both sides
4y=0
y=0 and x=-4
There you go. To see if these are correct plug the answers into both equation and it should come out good.
RJ Toftness
www.math-unlock.com